Polynomials: simplify
14t+14t+t=

Found 2 solutions by ewatrrr, lenny460:
Answer by ewatrrr(24785)

 
Hi,
14t+14t+t=
combining Like Terms
14t+14t+t= 29t

Answer by lenny460(1073)

14t + 14t + t

Combine like terms:

14t + 14t = 28t
28t + t = 29t

The Answer:

29t


obuong3@aol.com